A household consuming a bundle of goods X and Y such that MUX = 15 and MUY = 3 when the prices are such that pX = 5 and pY = 1 is maximizing its “utility”.
1
Expert's answer
2014-10-21T12:38:05-0400
MUx = 15, MUy = 3, Px = 5, Py = 1.
Household is maximizing its utility, if MUx/Px = MUy/Py, so:
15/5 = 3/1 = 3, so utility is maximized.